Circular No. 5660

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1992bk IN ESO 156-G8 M. Hamuy, Cerro Tololo Interamerican Observatory; and J. Maza,

University of Chile, report the discovery of a supernova (m_{pg} ∼ 18) by M. Wischnjewsky (University of Chile) on a 20-min unfiltered IIa-O plate taken by G. Valladares with the CTIO Curtis Schmidt telescope on Nov. 19.23 UT. SN 1992bk is located ∼ 12′′ east and 21′′ north of the nucleus of the elliptical galaxy ESO 156-G8 (α = 3^{h}41^{m}43^{s}.4, δ = −53^{o}46′55′′, equinox 1950.0); this galaxy is near the center of the rich cluster of galaxies CA 0340–538 (Sersic 1974, Ap. Space Sci. 28, 365). Confirmation of SN 1992bk was made by K.-P. Cheng (Goddard Space Flight Center) on CCD B and V images obtained on Nov. 21.25 with the CTIO 0.9-m telescope. A spectrogram (range 320.0-750.0 nm) was obtained by Maza on Nov. 21.2 with the CTIO 4-m telescope, revealing the characteristic Si II 635.5-nm feature of type-Ia supernovae. The minimum of the Si II absorption yields an expansion velocity of ∼ 10 700 km/s with respect to the galaxy’s rest frame. This spectrum shows resemblance to the spectrum of the type-Ia SN 1992A taken 7 days past maximum light. The spectrum of the nucleus of the parent galaxy yields a redshift of z = 0.058.

SUPERNOVA 1992bl IN ESO 291-G11 Hamuy and Maza further report the discovery of a supernova (m_{pg}

∼ 17) by R. Antezana (University of Chile), found on a 20-min unfiltered IIa-O Curtis Schmidt plate taken by Valladares on Nov. 20.07 UT. SN 1992bl is located ∼ 15′′ east and 22′′ south of the nucleus of the spiral galaxy ESO 291-G11 (α = 23^{h}12^{m}25^{s}.0, δ = −45^{o}00′36′′, equinox 1950.0). A spectrogram (taken as above by Maza) on Nov. 22.1 again shows the characteristic Si II 635.5-nm feature of type-Ia supernovae. The minimum of the Si II absorption yields an expansion velocity of ∼ 11 400 km/s with respect to the galaxy’s rest frame. This spectrum shows resemblance to that of SN 1992A taken at maximum light. The spectrum of the nucleus of the parent galaxy yields a redshift of z = 0.044. Confirmation of SN 1992bk was also made by Cheng (as above) on Nov. 23.06. Photoelectric photometry was obtained for this object by A. U. Landolt (Louisiana State University) with the CTIO 1-m telescope on Nov. 23.08, yielding V = 17.2, B–V = +0.07, U–B = –0.14.

1992 November 25 Daniel W. E. Green